eviant:

Math help pls Solve the equation:5+6x=7-3x

eviant:

@Vocaloid @dude

eviant:

Here's what I did:

dude:

Oh the t was unintentional? `5+6tx=7-3x`

eviant:

@dude yes,sorry

dude:

So you had the right idea of trying to remove one side but youre supposed to do the opposite sign to get rid of it (So instead of adding, you would subtract the 7)

1 attachment
eviant:

@dude

mhchen:

So you have \[-2 + 3x = 0\] so far right? Do you know what to do next?

eviant:

no

mhchen:

\[-2 + 3x = 0\] \[3x = 2\] Did you see what I did?

eviant:

how did -2 become positive?

mhchen:

\[-2 + 3x = 0\] Add 2 to both sides: \[-2 + 3x +2= 0+2\]

eviant:

Oh, so what do we do with 3x=2

mhchen:

We need to get x by itself, so do \[\frac{3x}{3} = \frac{2}{3}\]

eviant:

@mhchen 0.6?

mhchen:

Well you got \[x = \frac{2}{3}\] which is the same as 0.666666666666...... I think you can just write 2/3.